OpenStudy (anonymous):

How would i multipy this? Multiply. (2n + 2)(2n – 2) 4n2 – 4 4n2 – 4n – 4 4n2 + 2n – 4 4n2 + 4n – 4

jimthompson5910 (jim_thompson5910):

Similar Example: (5x + 7)(5x - 7) 5x(5x - 7) + 7(5x - 7) 25x^2 - 35x + 35x - 49 25x^2 - 49 So in the end, (5x + 7)(5x - 7) = 25x^2 - 49

OpenStudy (anonymous):

Try "FOIL" First. Outside. Inside. Last.

jimthompson5910 (jim_thompson5910):

A shortcut: (5x + 7)(5x - 7) = (5x)^2 - (7)^2 (5x + 7)(5x - 7) = 25x^2 - 49 This is using the difference of squares rule

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):